your investment has a 30% chance of losing 8%, a 40% chance of earning a 12% rate of return, and a 30% chance of earning a 15% return. What is your expected return in this investment.

A 8.8%

B 6.9%

C 7.14%

D 8.54%
Answer #1

A B 1 1 Prob. 30% 40% 30% 2 3 4 5 С Return -8% 12% 15% D Prob*return -2% 5% 5% 6.90% 2 3

Calc:

1 A B с D Prob. Return Prob* return 2 1 0.3 -0.08 =B2*C2 3 2 0.4 0.12 =B3*C3 4 3 0.3 0.15 =B4* C4 5 =SUM(D2:04)